OpenStudy (anonymous):

cos[tan^-1(-1) + cos^-1 (-4/5) How would I find the exact value of this one? Will FAN and MEDAL

OpenStudy (anonymous):

@zepdrix sorry to bother you again but can you help me again lol

zepdrix (zepdrix):

Figure this one out yet? :)

OpenStudy (anonymous):

Nope all I have is that tan inverse of -1 is -pi/4

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Ok that at least gets us started,\[\Large\rm \cos\left[-\frac{\pi}{4}+\arccos\left(-\frac{4}{5}\right)\right]\]Recall that the composition of a function and it's inverse just gives you the argument back. I'm saying this because... we're going to end up with cos(arccos(-4/5)) at some point, so we'll want to keep this in mind:\[\Large\rm \cos\left(\arccos\left(x\right)\right)=x\]

zepdrix (zepdrix):

Our inverse function represents some angle. So what we REALLY have, even though it might not look like it right now, is something like this: \(\Large\rm \cos(\theta+\beta)\)

zepdrix (zepdrix):

Remember your Cosine Angle Sum Identity?

OpenStudy (anonymous):

\[\cos(\theta+\beta)=\cos \theta \cos \beta - \sin \theta \sin \beta\]

OpenStudy (anonymous):

?

zepdrix (zepdrix):

Ya, we want to use that.

zepdrix (zepdrix):

Oh oh oh, we're going to end up with a sin(arccos) as well.. so I guess we better just do the triangle business ok ok ok

zepdrix (zepdrix):

So I'm saying that this arccos represents some arbitrary angle,\[\Large\rm \arccos\left(-\frac{4}{5}\right)=\theta\]We'll rewrite this relationship like this,\[\Large\rm \cos \theta=-\frac{4}{5}\]And draw a triangle to show the relationship.

zepdrix (zepdrix):

something like that, ya?

OpenStudy (anonymous):

Yep

zepdrix (zepdrix):

So using this information, \[\Large\rm \color{royalblue}{\arccos\left(-\frac{4}{5}\right)=\theta}\] Our problem,\[\Large\rm \cos\left[-\frac{\pi}{4}+\color{royalblue}{\arccos\left(-\frac{4}{5}\right)}\right]\]really looks like this,\[\Large\rm \cos\left[-\frac{\pi}{4}+\color{royalblue}{\theta}\right]\]

zepdrix (zepdrix):

Then we can apply our Angle Addition Formula for Cosine,\[\Large\rm =\cos\left(-\frac{\pi}{4}\right)\cos\left(\color{royalblue}{\theta}\right)-\sin\left(-\frac{\pi}{4}\right)\sin\left(\color{royalblue}{\theta}\right)\]

OpenStudy (anonymous):

Got it

zepdrix (zepdrix):

And then use the triangle to fill in the pieces related to theta, ya? :)

OpenStudy (anonymous):

Yes, should would that change sin theta

zepdrix (zepdrix):

yes, I guess if we look at our triangle, we get something like 3/5 for our sin(theta) ya?

OpenStudy (anonymous):

Yes

zepdrix (zepdrix):

\[\Large\rm =\cos\left(-\frac{\pi}{4}\right)\cos\left(\color{royalblue}{\theta}\right)-\sin\left(-\frac{\pi}{4}\right)\sin\left(\color{royalblue}{\theta}\right)\] \[\Large\rm =\cos\left(-\frac{\pi}{4}\right)\left(-\frac{4}{5}\right)-\sin\left(-\frac{\pi}{4}\right)\frac{3}{5}\]

zepdrix (zepdrix):

You'll have to simplify a little further :o Remember how to figure these out? Unit circle might help.

OpenStudy (anonymous):

Would the cos be 3pi/4's x

zepdrix (zepdrix):

hmm? 0_o

zepdrix (zepdrix):

we're taking the cosine of an angle, we don't end up with any pi business. we end up with one of those special values, ya? 1/2, sqrt(3)/2, sqrt(2)/2, 1, 0 one of those guys, remember?